OpenStudy (anonymous):

Find an exact value. cos 75°

OpenStudy (anonymous):

In radian? Or degree mode?

OpenStudy (anonymous):

@SGREBORN for example, like -sqrt(6)+sqrt(2)

OpenStudy (anonymous):

Is this topic on Imaginary numbers?

OpenStudy (raden):

first, setting 75 = 30+45 then use the formula : cos(A+B) = cosAcosB - sinAsinB here, given A=30 and B=45

OpenStudy (anonymous):

@RadEn so i got cos30cos45-sin30sin45 how do i simplify this?

OpenStudy (raden):

do u know what the values of cos30, cos 45, sin 30, and sin45 degrees ? they all the special degrees, u have to memorize it :)

OpenStudy (anonymous):

@RadEn ahh is this unit circle? I forgot:(

OpenStudy (raden):

it is trigonometry

OpenStudy (raden):

well, if u forgot i will refresh ur memories :) cos30 = 1/2 * sqrt(3) cos45 = 1/2 * sqrt(2) sin30 = 1/2 sin45 = 1/2 * sqrt(2) now, apply of them to formula above

OpenStudy (anonymous):

@RadEn so it will be {sqrt(6) - sqrt(2)}/4?

OpenStudy (raden):

yes, u are right

OpenStudy (anonymous):

@RadEn Thankyou!